OpenStudy (anonymous):

Determine whether the ratios 3/5 and 6/12 are proportional. a. no b. yes

OpenStudy (unklerhaukus):

if you convert those fractions to 60th's you can make a direct comparison

OpenStudy (anonymous):

im so confused

OpenStudy (unklerhaukus):

\[\frac35=\frac35\times\frac{12}{12}=\frac{3\times12}{5\times12}=\frac{??}{60}\]

OpenStudy (unklerhaukus):

\[\frac6{12}=\frac6{12}\times\frac55=\frac{?}{60}\]

OpenStudy (anonymous):

so no?